# On an Erdős-Szekeres Game

### v1

@@ 1 Introduction
In 2021, the current author proposed a variation on this game as part of a public lecture [4]. Instead of proposing a 2-player competitive game, the new version is collaborative, and instead of picking specific digits from a set, players append a new digit onto a permutation prefix. In particular, after the first $n$ moves, the current game state is a permutation $\pi\in\mathcal{S}_{n}$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move. After the first move, the game is always in state $\pi=1$. However, on subsequent moves, if a player chooses $m\in\left\{1,\dots,n+1\right\}$, we let $A=\left\{i\middle|\pi_{i}\leq m-1\right\}$ and let $B=\left\{i\middle|\pi_{i}\geq m\right\}$ and the new game state becomes $\pi^{\prime}=\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}m\in\mathcal{S}_{n+1}$ where $\pi_{i}^{\prime}=\pi_{i}$ if $i\in A$ and $\pi_{i}^{\prime}=\pi_{i}+1$ if $i\in B$. Notice that $\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}$ forms a $\pi$ pattern in $\pi^{\prime}$, that is, the relative order of the initial digits is unchanged. The game ends when the most recent move completes an $I_{a}$ pattern or a $J_{b}$ pattern. The collaborative version of this game was generated as a teaching tool to build intuition about Theorem 1 for audience members who were unfamiliar with the theorem. Optimal collaborative game play achieves a permutation of length $(a-1)(b-1)+1$.
@@
In this paper, we consider the 2-player competitive version of this game. In particular, as in the collaborative version, after the first $n$ moves, the current game state is a permutation of length $n$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move, appending it to the current permutation. We will primarily analyze the game where the first player who completes an $I_{a}$ pattern or a $J_{b}$ pattern loses. In contrast to the analysis of Harary, Sagan, and West’s game, where no general strategy was found, we give strategies that work for specific choices of $b$ but for any $a\geq b$.
@@ 2 Representation of Moves
We now make some observations about the shaded region of a board at any point in an $(a,b)$-game. First, by definition of eliminated cells, at any point in the game, the set of shaded and eliminated board cells forms one contiguous region. Moreover, this region tells us clearly which cells are available to be shaded by the next move of the game. We refer to cells that are not shaded and not eliminated as open cells.
@@
These four cases decribe the set of possible cells that can be shaded by various choices of $\pi_{n+1}$. We gave specific examples that showed $(1,d+1)$ and $(i+1,1)$ are possible. We also saw that each open cell that could be shaded by a choice of $\pi_{n+1}$ differs in row and/or column number by at most 1 from another open cell. In fact, the only time when both numbers change is when they are immediately below and immediately right of a shaded cell that forms a southeast corner of $S$. This uniquely describes the open cells that are edge adjacent to $S$.
@@
In the following sections, we articulate a strategy for winning this permutation game in terms of board shading.
@@ 7 Strategy for the misère game
However, if playing off the board results in a win, the penultimate move will be in column $a-1$ or row $b-1$, giving the player an opportunity to finish the appropriate monotone subsequence. Since playing in column $a-1$ or row $b-1$ gives the opponent a win, players seek to not play in this row and column. In other words, players wanting to win, try to restrict themselves to the $(b-2)\times(a-2)$ subboard. A player who claims cell $(a-2,b-2)$ then forces their opponent to play in the last row or last column, which allows the player who claimed $(a-2,b-2)$ to win. In other words, the misère game requires the same strategy as the original game, but on a board with one fewer row and one fewer column. This means that player 1 has a misère winning strategy when $4\leq b\leq 6$. The parity of $a$ determines the winner when $b=3$. And player 2 is guaranteed a misère win when $b=2$ merely by completing a decreasing $J_{2}$ pattern on their first turn.

### v2

@@ 1 Introduction
In 2009, Albert et. al. [1] considered the game of Harary, Sagan, and West as well as a number of generalizations. In particular, they also considered the play of the game on $\mathbb{Q}$ rather than on a finite set. In this latter setting they determined that the winner of the achievement game on $\mathbb{Q}$ with parameters $a$ and $b$ is the same as the winner of the avoidance game with parameters $a-1$ and $b-1$ (Proposition 7 of [1]). They also determined the winner of the achievement game for $a\geq b$ and $b\leq 5$ (Theorem 9 of [1]). The cases where $b=2$ and $b=3$ are simple, but they gave an explicit strategy for a first player win in the cases where $b=4$ and $b=5$. In addition, for arbitrarily large $b$, they determine that the game where $a=b$ is a first player win for $b\geq 4$.
In 2021, the current author proposed a variation on this game as part of a public lecture [5]. Instead of proposing a 2-player competitive game, the new version is collaborative, and instead of picking specific digits from a set, players append a new digit onto a permutation prefix. In particular, after the first $n$ moves, the current game state is a permutation $\pi\in\mathcal{S}_{n}$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move. After the first move, the game is always in state $\pi=1$. However, on subsequent moves, if a player chooses $m\in\left\{1,\dots,n+1\right\}$, we let $A=\left\{i\middle|\pi_{i}\leq m-1\right\}$ and let $B=\left\{i\middle|\pi_{i}\geq m\right\}$ and the new game state becomes $\pi^{\prime}=\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}m\in\mathcal{S}_{n+1}$ where $\pi_{i}^{\prime}=\pi_{i}$ if $i\in A$ and $\pi_{i}^{\prime}=\pi_{i}+1$ if $i\in B$. Notice that $\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}$ forms a $\pi$ pattern in $\pi^{\prime}$, that is, the relative order of the initial digits is unchanged. This relative-order requirement is equivalent to playing a collaborative form of the game of [3] on $\mathbb{Q}$ and is thus a variation on the competitive game of [1]. The game ends when the most recent move completes an $I_{a}$ pattern or a $J_{b}$ pattern. The collaborative version of this game was generated as a teaching tool to build intuition about Theorem 1 for audience members who were unfamiliar with the theorem. Optimal collaborative game play achieves a permutation of length $(a-1)(b-1)+1$.
@@
In this paper, we consider the 2-player competitive version of this game, which is a reformulation of the game on $\mathbb{Q}$ studied in [1]. In particular, as in the collaborative version, after the first $n$ moves, the current game state is a permutation of length $n$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move, appending it to the current permutation. We will primarily analyze the game where the first player who completes an $I_{a}$ pattern or a $J_{b}$ pattern loses. In contrast to the analysis of Harary, Sagan, and West’s game, where no general strategy was found, we give strategies that work for specific choices of $b$ but for any $a\geq b$. In contrast to the formulation in [1], we model the game as shading cells in a 2-dimensional grid while they model it by a building a ternary word with certain restrictions. For completeness, we show how to use the grid-shading model to see that the parity of $a$ determines the winner when $b=2$ and to give a strategy for a first player win in the avoidance game for $b=3$ and $b=4$. Since [1] focuses primarily on the attainment game, these strategies match their results for $b=3$, $b=4$, and $b=5$. In addition, we use our grid-shading model to give a winning strategy for the $b=5$ avoidance game, which goes beyond the scope of winning strategies given in [1].
@@ 2 Representation of Moves
We now make some observations about the shaded region of a board at any point in an $(a,b)$-game.
First, by definition of eliminated cells, at any point in the game, the set of shaded and eliminated board cells forms one contiguous region. Moreover, this region tells us clearly which cells are available to be shaded by the next move of the game. We refer to cells that are not shaded and not eliminated as open cells.
@@
These four cases describe the set of possible cells that can be shaded by various choices of $\pi_{n+1}$. We gave specific examples that showed $(1,d+1)$ and $(i+1,1)$ are possible. We also saw that each open cell that could be shaded by a choice of $\pi_{n+1}$ differs in row and/or column number by at most 1 from another open cell. In fact, the only time when both numbers change is when they are immediately below and immediately right of a shaded cell that forms a southeast corner of $S$. This uniquely describes the open cells that are edge adjacent to $S$.
@@
Finally, we remark that this region of shaded and eliminated cells encodes exactly the same information about $\pi$ as the ternary words used in [1]. The notation in Albert et. al. is motivated directly by Schensted’s bumping algorithm in [9] rather than Seidenberg’s proof. They represent game states as words on the symbols R, B, and P with at least one P, not starting with B, not ending with R, and not containing RB as a factor. Notice that the boundary of the shaded/eliminated region in our representation is contiguous. If we consider shaded and eliminated cells with bottom and/or right edges on the boundary of the shaded region, we can traverse these cells from lower left to upper right, and record R if only the bottom edge is on the boundary of the shaded region, B if only the right edge is on the boundary of the shaded region, and P if both the bottom and right edges are on the boundary of the shaded region. For example, the gray shaded and eliminated region corresponding to $\pi=163425$ shown in Figure 1 corresponds to the word $RPRPB$. The word notation of Albert et. al. has the advantage that it is language theoretic. The grid presentation in this paper has the advantage that one can play the game focused on geometric information, which can make it easier to quickly ascertain the state of the game without counting or computation.
In the following sections, we articulate a strategy for winning this permutation game in terms of board shading. The strategies for $2\leq b\leq 4$ exactly match the strategies that were presented in language theoretic terms in [1] while the strategy for $b=5$ is completely new and extends their results.
@@ 7 Strategy for the achievement game
However, if playing off the board results in a win, the penultimate move will be in column $a-1$ or row $b-1$, giving the player an opportunity to finish the appropriate monotone subsequence. Since playing in column $a-1$ or row $b-1$ gives the opponent a win, players seek to not play in this row and column. In other words, players wanting to win, try to restrict themselves to the $(b-2)\times(a-2)$ subboard. A player who claims cell $(a-2,b-2)$ then forces their opponent to play in the last row or last column, which allows the player who claimed $(a-2,b-2)$ to win. In other words, the achievement game requires the same strategy as the original game, but on a board with one fewer row and one fewer column. This means that player 1 has a winning strategy for the achievement game when $4\leq b\leq 6$. The parity of $a$ determines the winner when $b=3$. And player 2 is guaranteed a win for the achievement game when $b=2$ merely by completing a decreasing $J_{2}$ pattern on their first turn. This observation exactly matches Theorem 9 of [1].